OpenStudy (anonymous):

find domain f(x)= (2x+1) / (2x-1) and domain f(x)=(2x^2+5x+3) / (2x^2-5x-3)

OpenStudy (anonymous):

f(x) = (2x+1) / (2x-1) What values cannot be x.

OpenStudy (anonymous):

1 nd -1 ?

OpenStudy (kainui):

Literally what is the domain? The domain is just all the possible x-values you're allowed to pick, that's it! So if you look at the domain of: 1/x You'll see that you can pick any number EXCEPT for 0. Why? Well, you can't divide by zero, can you? So any time you have something at the bottom of a fraction with x's in it, set the bottom equal to zero and solve for what makes the bottom =0. Then you know you can't ever pick those numbers.

OpenStudy (anonymous):

but the bottom doesn't have any x's

OpenStudy (kainui):

Yes they both do.

OpenStudy (anonymous):

oh yea I see the bottom one the way it was typed I forgot it is on the bottom

OpenStudy (kainui):

\[ f(x)=\frac{ (2x+1) }{ (2x-1) }\] So when 2x-1=0 we know that will be dividing by zero. So we have to solve for x here to see when that is so we know that it's NOT in the domain.

OpenStudy (anonymous):

so we solve 2x-1=0?

OpenStudy (anonymous):

(-infty,1/2)u(1/2,infty)

OpenStudy (anonymous):

I got that for the bottom is that correct and do I need domain for both top and bottom or only bottom? Some1 plz help and I still need the second problem too